use the given conditions to write an equation for the line in point-slope form, slope-intercept for, and standard form. passing through (-6,-5) and parallel to the line whose equation is y-6=5/4(x-1).
Answer by Cromlix(4381)

Gradient = 5/4 (same gradient if parallel)
(-6, -5)
y - b = m(x - a)
y - (-5) = 5/4(x - (-6))
y + 5 = 5/4(x + 6)
y = 5/4x + 30/4 - 20/4
y = 5/4x + 10/4
Multiply thro' by 4
4y = 5x + 10
Hope this helps.
:-)